# PGA: Personalizing Grasping Agents with Single Human-Robot Interaction

### v1

@@ Abstract
Language-Conditioned Robotic Grasping (LCRG) aims to develop robots that ground and grasp objects based on natural language instructions. While robots capable of recognizing personal objects like “my wallet” can interact more naturally with non-expert users, current LCRG systems primarily limit robots to understanding only generic expressions. To this end, we introduce a task scenario GraspMine with a novel dataset that aims to locate and grasp personal objects given personal indicators via learning from a single human-robot interaction. To address GraspMine, we propose Personalized Grasping Agent (PGA), that learns personal objects by propagating user-given information through a Reminiscence—a collection of raw images from the user’s environment. Specifically, PGA acquires personal object information by a user presenting a personal object with its associated indicator, followed by PGA inspecting the object by rotating it. Based on the acquired information, PGA pseudo-labels objects in the Reminiscence by our proposed label propagation algorithm. Harnessing the information acquired from the interactions and the pseudo-labeled objects in the Reminiscence, PGA adapts the object grounding model to grasp personal objects. Experiments on GraspMine show that PGA significantly outperforms baseline methods both in offline and online settings, signifying its effectiveness and personalization applicability on real-world scenarios. Finally, qualitative analysis shows the effectiveness of PGA through a detailed investigation of results in each phase.
@@ I INTRODUCTION
Empowering robots with the ability to comprehend human natural language presents a formidable yet vital challenge within the realms of AI and Robotics [1]. This capability, which involves understanding and executing instructions couched in human language, allows for more intuitive human-robot interactions. Addressing this capability, the field has extensively studied Language-Conditioned Robotic Grasping (LCRG) [2, 3, 4], which focuses on robotic systems that ground and grasp objects based on language instructions.
Current approaches in LCRG predominantly rely on generic expressions when describing objects for manipulation, leading to less intuitive human-robot communication. For instance, a person might instinctively instruct, “Get my AirPods”. However, with current robotic systems that operate on generic understandings, one might need to instruct, “Get the white pebble-shaped object in front of the blue cup”. This forces users to craft their instructions to suit the robot’s limited knowledge base, what can be termed as robot-centric directives. However, many non-expert users find this robot-centric directives both unfamiliar and cumbersome. To truly facilitate more intuitive user-centric interactions, robots should have more shared personal knowledge with their users beyond the generic, understanding personalized expressions.
To address this problem, we introduce a novel personalized task scenario of LCRG called GraspMine with a benchmark dataset. GraspMine aims to locate and grasp personal objects given a personal indicator, e.g., “my sleeping pills”, that robotic systems with generic knowledge may not handle properly. According to the field of Personalization [5, 6], personalizing with a few examples is a common scenario since collecting a lot of personalized data for each user is infeasible in real-world applications. Considering this aspect, GraspMine requires robots to learn personal objects via minimal human-robot interaction, i.e., a one-time verbal introduction of a personal object, as depicted in Fig. 1. In GraspMine dataset, we consider learning of 96 personal objects, each with its own indicator. GraspMine contains five distinct test splits, each sample including an image containing multiple objects, a personal indicator, and the associated object’s coordinates.
@@
Figure 1: Personalizing grasping agents with single human-robot interaction. Upon the user’s introduction of a personal object, it retrieves the identical objects from its visual reminiscence. Leveraging the retrieved objects, the robot subsequently engages in an integrated learning process of the personal object. The personalized grasping agent can finally comprehend the personalized expression, grasping the personal object.
@@
To address GraspMine, we propose Personalized Grasping Agent (PGA), that learns personal objects by propagating the information acquired from a user through a Reminiscence. Specifically, PGA first constructs the Reminiscence—a collection of raw images from the user’s environment. Then PGA acquires personal object information in two successive steps: human-robot interaction and robot-object interaction. In human-robot interaction, a user introduces a personal item to the robot with a personal indicator, e.g.,my sleeping pills”. In robot-object interaction, the robot inspects the object from multiple perspectives, obtaining diverse images of personal objects. Leveraging the acquired information of the personal object, PGA pseudo-labels objects in the Reminiscence by propagating personal indicators based on the visual features of the objects, which stems from label propagation algorithm [7]. We term this Propagation through Reminiscence. Harnessing the data obtained from the interactions and the pseudo-labeled objects in Reminiscence, PGA adapts the object grounding model, and with the personalized model, it adeptly grasps the user-specified personal objects.
To validate the effectiveness of PGA, we conduct both offline (i.e., object discovery) and online (i.e., real-world LCRG) experiments on GraspMine. In offline experiments, PGA outperforms several baseline methods, and even shows comparable results with a fully-supervised method that employs extensive annotations within the Reminiscence. Moreover, PGA also robustly performs in an online setting, demonstrating its real-world personalization applicability. Concluding our analysis qualitatively, we present visual examples from various phases of PGA, offering insights into the mechanisms that contribute to its robust performance.
@@ II RELATED WORK
Language-Conditioned Robotic Grasping (LCRG) focus on the robot’s capability to grasp objects based on human instructions given in natural language. Numerous studies [8, 9, 4, 10, 11, 12, 3, 13] have employed fully supervised training approaches to teach robots how to locate objects based on natural language instructions, making use of public datasets [14, 15, 16]. However, these supervised learning methods reach their limit when robots have to recognize or interact with personal objects absent during the training. A study [17] curated its own dataset comprising unique objects. Still, it depended on a supervised learning, which necessitates exhaustive annotations of language instructions. GVCCI [2] introduced a paradigm shift by unveiling an unsupervised approach to adapt to the user’s personal objects. While GVCCI enabled robots to automatically learn the visual nuances of personal objects, it remained constrained. The learning was heavily reliant on basic-level categories [18] and attributes inferred by pretrained classifiers [19, 20]. A major limitation of such methodologies is their grounding of visual objects in basic-level categories. This forces non-expert users to use robot-centric instructions, a pain point discussed in Sec. I. Some studies [21, 22] delved deeper, augmenting robots’ knowledge with semantic or intended meanings of objects, typified by phrases like “I am thirsty”. While these studies enable robots to grasp user intentions, they also rely on fully-supervised learning, necessitating extensive annotations. Unlike previous works, PGA can efficiently learn the user’s personal objects and their corresponding language descriptors with just a single user-robot interaction.
@@
Personalization has become an important factor within various domains of Machine Learning [23, 24, 25, 26, 27, 28, 29, 30]. Personalization has also been studied in the field of Robotics, such as dressing assistance [31, 32, 33] and tidy-up task [34]. While these works of Personalization focus on the user’s personal preferences, our work focus on transferring the user’s personal knowledge, standing as a pioneering work in the realm of personalized LCRG. Two studies [5, 6] align closely with our research objectives, focusing on learning visual and linguistic representations of personal objects. The work [6] utilizes personal videos paired with transcripts to learn personal objects and their corresponding names, without the need for explicit human labeling. Yet, their focus lies on temporal grounding in videos, overlooking the crucial aspect of spatial grounding. This limitation renders their approach unsuitable for LCRG, where pinpointing the precise object location is imperative. The work [5] broadens this focus to include spatial grounding. However, they assume that fine-grained annotations (e.g., segmentation masks) and personal object names are readily available. In contrast, PGA only requires a single human-robot interaction without any further annotations. So, non-expert users can easily be involved in collecting the training data due to the intuitive and streamlined interface.
@@ III METHOD
Figure 2: Overview of Personalized Grasping Agent (PGA). (a) Initially, PGA gathers a collection of raw images, termed as Reminiscence, from the user’s environment. With the personal indicator provided by the user from (b), unlabelled objects in the Reminiscence are pseudo-labelled via (c) Propagation through Reminiscence. It’s vital to note that certain objects, particularly those not introduced by the user (e.g., non-personal objects), remain unlabelled by our algorithm. Ultimately, PGA employs all the object nodes with labels (colored nodes) to train the Personalized Object Grounding Model.
@@ III-A Reminiscence Construction
Personalized Grasping Agent (PGA) initiates by constructing the Reminiscence, as illustrated in Fig. 2-(a). PGA first collects a set of $N$ images from the user’s environment, represented as =$\mathcal{R}=\{I^{\mathcal{R}}_{n}\}_{n=1}^{N}$, which we term as Reminiscence. PGA detects all the objects $\{o^{\mathcal{R}}_{m}\}_{m=1}^{M}$ from $\mathcal{R}$, via an off-the-shelf object detector [20], and constructs a node for each detected objects, embedding them with the pretrained visual encoder $f(\cdot)$, DINO [35]. These object nodes, represented as $O^{\mathcal{R}}=\{\mathbf{o}^{\mathcal{R}}_{m}\}_{m=1}^{M}=\{f(o^{\mathcal{R}}_{m})\}_{m=1}^{M}$ where $M\gg N$, are tagged with the meta-information – an image from $\mathcal{R}$ and the bounding box (bbox) coordinates of the detected object – which will be leveraged in subsequent processes. Note that nodes in $O^{\mathcal{R}}$ are without the object’s personal indicators.
@@ III-B Object Information Acquisition
To acquire the information of personal objects, PGA goes through two successive steps as shown in Fig. 2-(b): human-robot interaction and robot-object interaction. In human-robot interaction, the user introduces their personal object $o$ by displaying the object and verbally describing it. The description consists of two indicators: a general indicator $u^{G}_{o}$, e.g., “the object in front”, and a personal indicator $u_{o}$, e.g., “my sleeping pills”. Using its visual perception $\mathcal{I}_{o}$ and $u^{G}_{o}$, PGA predicts the bbox coordinates $b_{o}$ of the object via GVCCI [2]. PGA then constructs an object node $\mathbf{o}=f(o)$ using pretrained visual encoder $f(\cdot)$ [35]. This node is tagged with the meta-information including $\mathcal{I}_{o}$, $b_{o}$, and $u_{o}$.
In robot-object interaction, PGA subsequently examines the personal object $o$ from multiple views, aiming to capture its distinctive visual appearance. By leveraging $b_{o}$, PGA grasps and rotates the object capturing $A$ images $\{\mathcal{I}_{o,a}\}_{a=1}^{A}$ for each object, each image offering a distinct view of the object (see the bottom right of Fig. 2). As in the earlier step, PGA predicts an object bbox coordinates, $\{b_{o,a}\}_{a=1}^{A}$ of a set of objects $\{{o_{a}}\}_{a=1}^{A}$. PGA constructs a set of object nodes $O^{view}=\{{\mathbf{o}_{a}}\}_{a=1}^{A}=\{{f(o_{a})}\}_{a=1}^{A}$, where each node $\mathbf{o}_{a}$ is tagged with the meta-information $\mathcal{I}_{o,a}$, $b_{o,a}$, and $u_{o}$. Finally, each personal object is associated with $\mathbf{o}$ and $O^{view}$, i.e., $A+1$ nodes that are tagged with personal indicators. Across all personal objects, the aggregation of nodes with personal indicators forms $O^{\mathcal{K}}$, while collective personal indicators form $U$, as described in lines 1-6 of Algorithm 1.
@@ III-C Propagation through Reminiscence
Propagation through Reminiscence draws inspiration from the semi-supervised method of label propagation [7, 36], as detailed in Algorithm 1 and depicted in Fig. 2-(c). The goal is to pseudo-label (tag) the missing indicators in $O^{\mathcal{R}}$ by propagating personal indicators from $O^{\mathcal{K}}$. Specifically, for every node $\mathbf{o}^{\mathcal{R}}\in O^{\mathcal{R}}$, PGA computes an affinity score, $S(\mathbf{o}^{\mathcal{R}},u)$, for each possible personal indicator $u\in U$. This score is the average cosine similarity $\phi$ between node $\mathbf{o}^{\mathcal{R}}$ and every known object node $\mathbf{o}^{\mathcal{K}}$ tagged with indicator $u$;
@@
$S(\mathbf{o}^{\mathcal{R}},u)=\frac{1}{N(u)}\sum_{\mathbf{o}^{\mathcal{K}}\in O^{\mathcal{K}}}I(u_{\mathbf{o}^{\mathcal{K}}}=u)\cdot\phi(\mathbf{o}^{\mathcal{R}},\mathbf{o}^{\mathcal{K}});$ (2)
where $N(u)$ denotes the number of $\mathbf{o}^{\mathcal{K}}$s tagged with indicator $u$ and $I(\cdot)$ is the indicator function that takes the value of 1 if the condition is true, and 0 otherwise. Note that $u_{\mathbf{o}^{\mathcal{K}}},u\in U$. Based on these scores, PGA assigns the indicator $u$ with the highest affinity score to $\mathbf{o}^{\mathcal{R}}$; $u_{o^{\mathcal{R}}}={\text{argmax}}\{S(\mathbf{o}^{\mathcal{R}},u)~{}|~{}u\in U\}$; if the highest score is above the certain threshold. After pseudo-labeling $\mathbf{o}^{\mathcal{R}}$ with the most likely personal indicator, $O^{\mathcal{K}}$ is updated to incorporate the propagated node $\mathbf{o}^{\mathcal{R}}$, and this updating process is iteratively performed until the proportion of altered nodes is less than 10% – until saturated. Propagation through Reminiscence allows PGA to effectively gather pseudo-labeled examples, thereby fostering robust, data-rich learning of the user’s personal object.
Algorithm 1 Propagation through Reminiscence
@@
19:         $\forall u\in U$, $S(\mathbf{o}^{\mathcal{R}},u)\leftarrow S_{u}/N_{u}$

### v2

@@ Abstract
Language-Conditioned Robotic Grasping (LCRG) aims to develop robots that comprehend and grasp objects based on natural language instructions. While the ability to understand personal objects like my wallet facilitates more natural interaction with human users, current LCRG systems only allow generic language instructions, e.g., the black-colored wallet next to the laptop. To this end, we introduce a task scenario GraspMine alongside a novel dataset aimed at pinpointing and grasping personal objects given personal indicators via learning from a single human-robot interaction, rather than a large labeled dataset. Our proposed method, Personalized Grasping Agent (PGA), addresses GraspMine by leveraging the unlabeled image data of the user’s environment, called Reminiscence. Specifically, PGA acquires personal object information by a user presenting a personal object with its associated indicator, followed by PGA inspecting the object by rotating it. Based on the acquired information, PGA pseudo-labels objects in the Reminiscence by our proposed label propagation algorithm. Harnessing the information acquired from the interactions and the pseudo-labeled objects in the Reminiscence, PGA adapts the object grounding model to grasp personal objects. This results in significant efficiency while previous LCRG systems rely on resource-intensive human annotations—necessitating hundreds of labeled data to learn my wallet. Moreover, PGA outperforms baseline methods across all metrics and even shows comparable performance compared to the fully-supervised method, which learns from 9k annotated data samples. We further validate PGA’s real-world applicability by employing a physical robot to execute GrsapMine. Code and data are publicly available at https://github.com/JHKim-snu/PGA.
@@ I INTRODUCTION
Empowering robots with the ability to comprehend human natural language presents a formidable yet vital challenge within the realms of AI and Robotics [1]. This capability, which involves understanding and executing human language instructions, allows for more intuitive human-robot interactions. The researchers have studied such capability in the context of Language-Conditioned Robotic Grasping (LCRG) [2, 3, 4, 5, 6, 7, 8, 9, 10, 11], which focuses on robotic systems that ground and grasp objects based on language instructions.
@@
Figure 1: Personalizing grasping agents with single human-robot interaction. Upon the user’s introduction of a personal object, it retrieves the identical objects from its visual reminiscence. Leveraging the retrieved objects, the robot subsequently engages in an integrated learning process of the personal object. The Personalized Grasping Agent (PGA) can finally comprehend and grasp the personal object.
Current approaches in LCRG [3, 4, 5, 6, 7, 8, 9, 10, 11, 2] predominantly rely on generic language expressions when describing objects for manipulation, leading to less intuitive human-robot communication. For instance, a person might instinctively instruct, “Get my wallet”. However, with current LCRG systems that operate on generic instructions, one might need to instruct, “Get the black-colored wallet next to the laptop”. This forces users to craft their instructions to suit the robot’s limited knowledge base, what can be termed as robot-centric directives. However, many non-expert users find this robot-centric directives both unfamiliar and cumbersome. To truly facilitate more intuitive user-centric interactions, robots should have more shared personal knowledge with their users beyond the generic, understanding personal objects.
To address this problem, we introduce a novel personalized task scenario of LCRG called GraspMine with a benchmark dataset. GraspMine aims to locate and grasp personal objects given a personal indicator, e.g., “my sleeping pills”, that robotic systems with generic knowledge may not handle properly. According to the field of Personalization [12, 13], personalizing with a few examples is a common scenario since collecting a lot of personalized data for each user is infeasible in real-world applications. Considering this aspect, GraspMine requires robots to learn personal objects via minimal human-robot interaction, i.e., a one-time verbal introduction of a personal object, as depicted in Fig. 1. In GraspMine dataset, we consider learning of 96 personal objects, containing five distinct test splits (Heterogeneous, Homogeneous, Paraphrased, and Cluttered), each sample including an image containing multiple objects, a personal indicator, and the associated object’s coordinates.
@@
However, with the prevailing learning strategies of current LCRG systems [3, 4, 5, 6, 7, 8, 9, 10, 11], supervised learning that requires vast amounts of data, it is challenging to tackle GraspMine with only a single sample. To address GraspMine, we propose Personalized Grasping Agent (PGA), that learns personal objects by propagating the information acquired from a user through a Reminiscence. Specifically, PGA first constructs the Reminiscence—a collection of raw images from the user’s environment. Then PGA acquires personal object information in two successive steps: human-robot interaction and robot-object interaction. In human-robot interaction, a user introduces a personal item to the robot with a personal indicator, e.g.,my sleeping pills”. In robot-object interaction, the robot inspects the object from multiple perspectives, obtaining diverse images of personal objects. Leveraging the acquired information of the personal object, PGA pseudo-labels objects in the Reminiscence by propagating personal indicators based on the visual features of the objects, inspired from the label propagation algorithm [14]. We term this Propagation through Reminiscence. Harnessing the data obtained from the interactions and the pseudo-labeled objects in Reminiscence, PGA adapts the object grounding model, and with the personalized model, it adeptly grasps the user-specified personal objects.
In the experiment, we provide baselines for GraspMine, including Direct and Supervised methods. Direct employs supervised learning with only one sample per object acquired from human-robot interaction, while Supervised represents an upper-bound model trained with nearly 9k annotated samples. In offline experiments focusing on object grounding, PGA outperforms the Direct model by an average absolute increment of approximately 30%, and even demonstrates comparable results to the Supervised method. Notably, we observe performance improvements in offline experiments as the number of images in Reminiscence increases. Transitioning to the online setting, we validate PGA’s real-world applicability by employing a physical robot to execute GraspMine. To qualitatively conclude our analysis, we present visual examples from different phases of PGA, providing insights into its robust performance.
To summarize, our contributions are mainly three-fold:
  • We introduce a novel personalized task scenario in LCRG called GraspMine, aimed at grounding and grasping personal objects based on personal indicators. This scenario fills a gap in current LCRG systems, which primarily rely on generic expressions, thus enhancing the intuitive nature of human-robot interactions.
  • We propose Personalized Grasping Agent (PGA) as a strong baseline for GraspMine. PGA learns to ground personal objects by leveraging information acquired from a single human-robot interaction, robot-object interaction, and Propagation through Reminiscence.
  • We provide comprehensive experimental validation of PGA’s performance against baselines, showcasing its effectiveness in grounding personal objects. Additionally, we demonstrate the PGA’s practical applicability through deployment on a real-world physical robot.
@@ II-A Language-Conditioned Robotic Grasping (LCRG)
Language-Conditioned Robotic Grasping (LCRG) focuses on the robot’s capability to grasp objects based on human instructions given in natural language. Numerous studies [3, 4, 5, 6, 7, 8, 9, 10] have employed fully supervised training approaches to teach robots how to locate objects based on natural language instructions, making use of public datasets [15, 16, 17]. However, these supervised learning methods reach their limits when robots have to recognize or interact with personal objects unseen during the training. A study [11] curated its own dataset comprising unique objects. Still, it depended on supervised learning, which necessitates exhaustive annotations of language instructions. GVCCI [2] introduced a paradigm shift by unveiling an unsupervised approach to adapt to the user’s personal objects. While GVCCI enabled robots to automatically learn the visual nuances of personal objects, it remained constrained. The learning was heavily reliant on generic object categories [18] and attributes inferred by pretrained classifier [19]. This forces non-expert users to use robot-centric instructions, a pain point discussed in Sec. I. Some studies [20, 21] delved deeper, augmenting robots’ knowledge with semantic or intended meanings of objects, typified by phrases like “I am thirsty”. While these studies enable robots to grasp user intentions, they also rely on fully-supervised learning that can not be applied to GraspMine, necessitating extensive annotations. Unlike previous works, PGA can efficiently learn the user’s personal objects and their corresponding language descriptors with just a single user-robot interaction.
II-B Personalization
@@
Personalization has become an important factor within various domains of Machine Learning [22, 23, 24, 25, 26, 27, 28, 29]. Personalization has also been studied in the field of Robotics, such as dressing assistance [30, 31, 32] and tidy-up task [33]. While these works of Personalization focus on the user’s personal preferences, our work focuses on transferring the user’s personal knowledge, standing as a pioneering work in the realm of personalized LCRG. Two studies [12, 13] align closely with our research objectives, focusing on learning visual and linguistic representations of personal objects. The work [13] utilizes personal videos paired with transcripts to learn personal objects and their corresponding names, without the need for explicit human labeling. Yet, their focus lies on temporal grounding in videos, overlooking the crucial aspect of spatial grounding. This limitation renders their approach unsuitable for LCRG, where pinpointing the precise object location is imperative. The work [12] broadens this focus to include spatial grounding. However, they assume that fine-grained annotations (e.g., segmentation masks) and personal object names are readily available. In contrast, PGA only requires a single human-robot interaction without any further annotations. So, non-expert users can easily be involved in collecting the training data due to the intuitive and streamlined interface.
Figure 2: Overview of Personalized Grasping Agent (PGA). (a) Initially, PGA gathers a collection of raw images, termed as Reminiscence, from the user’s environment. With the personal indicator provided by the user from (b), unlabeled objects in the Reminiscence are pseudo-labeled via (c) Propagation through Reminiscence. It’s vital to note that certain objects, particularly those not introduced by the user (e.g., non-personal objects), remain unlabeled by our algorithm. Ultimately, PGA employs all the object nodes with labels (colored nodes) to train the Personalized Object Grounding Model.
@@ III-A Reminiscence Construction
Personalized Grasping Agent (PGA) initiates by constructing the Reminiscence, as illustrated in Fig. 2-(a). PGA first collects a set of $N$ images from the user’s environment, represented as =$\mathcal{R}=\{I^{\mathcal{R}}_{n}\}_{n=1}^{N}$, which we term as Reminiscence. PGA detects all the objects $\{o^{\mathcal{R}}_{m}\}_{m=1}^{M}$ from $\mathcal{R}$, via an off-the-shelf object detector [19], and constructs a node for each detected objects, embedding them with the pretrained visual encoder $f(\cdot)$, DINO [34]. These object nodes, represented as a set of vector embeddings $O^{\mathcal{R}}=\{\mathbf{o}^{\mathcal{R}}_{m}\}_{m=1}^{M}=\{f(o^{\mathcal{R}}_{m})\}_{m=1}^{M}$ where $M\gg N$, are tagged with the meta-information – an image from $\mathcal{R}$ and the bounding box (bbox) coordinates of the detected object – which will be leveraged in subsequent processes. Note that nodes in $O^{\mathcal{R}}$ are without the object’s personal indicators.
@@ III-B Object Information Acquisition
To acquire the information of personal objects, PGA goes through two successive steps as shown in Fig. 2-(b): human-robot interaction and robot-object interaction. In human-robot interaction, the user introduces their personal object $o$ by displaying the object and verbally describing it. The description consists of two indicators: a general indicator $u^{G}_{o}$, e.g., “the object in front”, and a personal indicator $u_{o}$, e.g., “my sleeping pills”. Using its visual perception $I_{o}$ and $u^{G}_{o}$, PGA predicts the bbox coordinates $b_{o}$ of the object via GVCCI [2]. PGA then constructs an object node $\mathbf{o}=f(o)$ using pretrained visual encoder $f(\cdot)$ [34]. This node is tagged with the meta-information including $I_{o}$, $b_{o}$, and $u_{o}$.
In robot-object interaction, PGA subsequently examines the personal object $o$ from multiple views, aiming to capture its distinctive visual appearance. By leveraging $b_{o}$, PGA grasps and rotates the object capturing $A$ images $\{I_{o,a}\}_{a=1}^{A}$ for each object, each image offering a distinct view of the object (see the bottom right of Fig. 2). PGA extracts a set of objects $\{{o_{a}}\}_{a=1}^{A}$ and their bounding box coordinates $\{b_{o,a}\}_{a=1}^{A}$ from the captured images $\{I_{o,a}\}_{a=1}^{A}$. PGA constructs a set of object nodes $O^{view}=\{{\mathbf{o}_{a}}\}_{a=1}^{A}=\{{f(o_{a})}\}_{a=1}^{A}$, where each node $\mathbf{o}_{a}$ is tagged with the meta-information $I_{o,a}$, $b_{o,a}$, and $u_{o}$. Finally, each personal object is associated with $\mathbf{o}$ and $O^{view}$, i.e., $A+1$ nodes that are tagged with personal indicators. Across all personal objects, the aggregation of nodes with personal indicators forms $O^{\mathcal{K}}$, while collective personal indicators form $U$, as described in lines 1-6 of Algorithm 1.
@@ III-C Propagation through Reminiscence
Propagation through Reminiscence draws inspiration from the semi-supervised method of label propagation [14, 35], as detailed in lines 7-20 of Algorithm 1 and Fig. 2-(c). The goal is to pseudo-label (tag) the missing indicators in $O^{\mathcal{R}}$ by propagating personal indicators from $O^{\mathcal{K}}$. For every node $\mathbf{o}^{\mathcal{R}}\in O^{\mathcal{R}}$, PGA computes an affinity score, $S(\mathbf{o}^{\mathcal{R}},u)$, for each possible personal indicator $u\in U$. This score is the average cosine similarity $\phi$ between node $\mathbf{o}^{\mathcal{R}}$ and every known object node $\mathbf{o}^{\mathcal{K}}$ tagged with indicator $u$;
@@
$S(\mathbf{o}^{\mathcal{R}},u)=\frac{1}{N(u)}\sum_{\mathbf{o}^{\mathcal{K}}\in O^{\mathcal{K}}}\mathbbm{1}(u_{\mathbf{o}^{\mathcal{K}}}=u)\cdot\phi(\mathbf{o}^{\mathcal{R}},\mathbf{o}^{\mathcal{K}});$ (2)
where $N(u)$ denotes the number of $\mathbf{o}^{\mathcal{K}}$s tagged with indicator $u$ and $\mathbbm{1}(\cdot)$ is the indicator function that takes the value of 1 if the condition is true, and 0 otherwise. Note that $u_{\mathbf{o}^{\mathcal{K}}},u\in U$. Based on these scores, PGA assigns the indicator $u$ with the highest affinity score to $\mathbf{o}^{\mathcal{R}}$; $u_{o^{\mathcal{R}}}={\text{argmax}}\{S(\mathbf{o}^{\mathcal{R}},u)~{}|~{}u\in U\}$; if the highest score is above the certain threshold. After pseudo-labeling $\mathbf{o}^{\mathcal{R}}$ with the most likely personal indicator, $O^{\mathcal{K}}$ is updated to incorporate the propagated node $\mathbf{o}^{\mathcal{R}}$, and label propagation repeats until the percentage of relabeled nodes is less than 10%. Propagation through Reminiscence allows PGA to automatically gather pseudo-labeled examples, alleviating the scarcity of labeled data in GraspMine.
Algorithm 1 Personalized Grasping Agent (PGA)
@@
19:         $\forall u\in U$, $S(\mathbf{o}^{\mathcal{R}},u)\triangleq S_{u}/N_{u}$